a total of $54,000 is to be invested, some in bonds and some in certificates of deposit (CDs). If the amount invested in bonds is to exceed that in CDs by $4,000, how much will be invested in each type of investment?
Answer by richwmiller(17219)

b+c=54000
b=c+4000
b-c=4000
b+c=54000
add
2b=58000
b=29000
c=25000